OpenStudy (mackenzie17):

Consider the functions f(x) = x2 − 13 and g(x) = x + 5. What is the value of f[g(−4)]? −12 3 −13 8

OpenStudy (evoker):

plug -4 into the g function and then plug that answer into the f function.

OpenStudy (mackenzie17):

thank you

OpenStudy (mackenzie17):

wait would the g function be -4-5 and then i work it out

OpenStudy (mackenzie17):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

it would be +5 (not -5)

jimthompson5910 (jim_thompson5910):

g(x) = x + 5 g(-4) = -4 + 5 ... replace every x with -4 g(-4) = ________ fill in the blank and tell me what you get

OpenStudy (mackenzie17):

-1

OpenStudy (mackenzie17):

no 4 sorry

jimthompson5910 (jim_thompson5910):

-1 is close but not quite there

jimthompson5910 (jim_thompson5910):

4 isn't correct either

OpenStudy (mackenzie17):

1

jimthompson5910 (jim_thompson5910):

draw a number line. Start at -4 and move 5 spots to the right

jimthompson5910 (jim_thompson5910):

yes +1 or just 1

jimthompson5910 (jim_thompson5910):

g(-4) = 1 this means that g(-4) is the same as 1

jimthompson5910 (jim_thompson5910):

f[g(−4)] would turn into f(1) which is equal to what?

OpenStudy (mackenzie17):

is it -11

jimthompson5910 (jim_thompson5910):

f(x) = x^2 - 13 f(1) = (1)^2 - 13 ... replace every x with 1 f(1) = ________ fill in the blank and tell me what you get

jimthompson5910 (jim_thompson5910):

`is it -11` the answer isn't -11

jimthompson5910 (jim_thompson5910):

you're close though

OpenStudy (mackenzie17):

11

jimthompson5910 (jim_thompson5910):

what is (1)^2 equal to?

OpenStudy (mackenzie17):

2

jimthompson5910 (jim_thompson5910):

1^2 is not the same as 1 times 2

jimthompson5910 (jim_thompson5910):

if you wanted to say 1 times 2, you'd write 1*2

jimthompson5910 (jim_thompson5910):

https://www.mathsisfun.com/exponent.html 1^2 means 1 with the exponent of 2 1^2 is the same as \(\Large 1^2\)

OpenStudy (mackenzie17):

ok so it it 12

jimthompson5910 (jim_thompson5910):

closer but no

jimthompson5910 (jim_thompson5910):

Start at 1. Then move 13 units to the left on the number line

OpenStudy (mackenzie17):

-13

jimthompson5910 (jim_thompson5910):

let's say you started at 1 and moved 3 spots to the left. Where do you end up?

OpenStudy (mackenzie17):

-3

jimthompson5910 (jim_thompson5910):

look at a numberline like this http://www.mathatube.com/sitebuilder/images/number-line-600x271.png

OpenStudy (mackenzie17):

-2

jimthompson5910 (jim_thompson5910):

yes now do the same for 1-13 start at 1 and move 13 units to the left

OpenStudy (mackenzie17):

-12

jimthompson5910 (jim_thompson5910):

yes so 1-13 is equal to -12

jimthompson5910 (jim_thompson5910):

meaning that -12 is the answer

OpenStudy (mackenzie17):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome